OpenStudy (anonymous):

Chad drove 168 miles in 3 hours. How many miles per hour did Chad drive? A. 171 B. 504 C. 65 D. 56

OpenStudy (anonymous):

@Daniel56k

OpenStudy (anonymous):

@K12IsAwesome @usercode3rror

OpenStudy (anonymous):

helppp

OpenStudy (anonymous):

@ktbug3 @YUMMY123

OpenStudy (anonymous):

@why3

OpenStudy (anonymous):

what do you think

OpenStudy (anonymous):

will medal

OpenStudy (anonymous):

D

OpenStudy (anonymous):

@Tboi

OpenStudy (anonymous):

Factor out 3 from the numerator and the denominator $168=3\cdot\:56\mathrm{,\:\quad}3=3\cdot\:1$168=3· 56, 3=3· 1 $=\frac{3\cdot\:56}{3\cdot\:1}$=3· 563· 1 $=56$=56

OpenStudy (anonymous):

there you go

OpenStudy (anonymous):

I think D

OpenStudy (anonymous):

168/3 is 56

OpenStudy (anonymous):

its 56

OpenStudy (anonymous):

=56

OpenStudy (anonymous):

so i ma right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thxs can u help again

OpenStudy (anonymous):

$=\frac{3\cdot\:56}{3\cdot\:1}$=3· 563· 1 $=56$=56

OpenStudy (anonymous):

i will!!

OpenStudy (anonymous):

Alice and Brian have cookies in a ratio of 4:3. Brian and Cori have cookies in a ratio of 6:4. What is the ratio of Alice’s cookies to Cori’s cookies? A. 8 to 4 B. 6 to 4 C. 3 to 6 D. 4 to 8 i think C

OpenStudy (anonymous):

medal

OpenStudy (anonymous):

okay gimme a sec

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

i think you are right! C!

OpenStudy (anonymous):

your right

OpenStudy (anonymous):

yay thxs

OpenStudy (anonymous):

any more